# Superfluid Dark Matter

### v1

@@ 2.6.1 Quasi-particle description
From the first term,$-4\pi G\rho$, we see that long-wave phonons are now unstable. This instability bears resemblance to the familiar Jeans instability in fluids under the influence of gravity [165, 166]. In other words, large-scale fluctuations of the condensate grow due to gravity, which cannot be counterbalanced by repulsive self-interactions on those scales. See also Ref. [167] for a generalization on a cosmological background.
@@ 3 Superfluid dark matter in galaxies
As shown above, the effect of gravity is balanced by two contributions. The first is the gradient of the polytropic equation of state, $-\frac{1}{\rho}\vec{\nabla}P$, which encodes the role of self-interactions in the evolution of the system. The second is the quantum pressure contribution, $\frac{1}{2m^{2}}\vec{\nabla}\left(\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}\right)$, which arises from the wave-like nature of the light bosonic particles. Depending on which dominates at the characteristic scales of interest—typically $\sim{\rm kpc}$ scales in astrophysical contexts—different dynamical regimes emerge. If quantum pressure is the dominant contribution, the system behaves according to the fuzzy DM model. In the opposite limit, where self-interactions dominate, the dynamics are governed by the SDM model, which is the primary focus of this review.
@@ 3.1 Hydrodynamics of the dark matter superfluid
The first model of SDM was proposed by J. Goodman in the seminal paper [6]. In this work, the author introduced DM as a scalar boson with repulsive two-body interactions, which is degenerate at typical galactic densities and interacts through repulsive short-range interactions. The important implication of the original model is that the ground state pressure sourced by self-interactions can sustain self-gravitating, zero-temperature configurations. We refer to these structures as non-topological solitons. This stands in contrast to self-interacting DM models, where cores are prone to gravothermal collapse upon cooling [78, 169], which was one of the original motivations behind the introduction of the model.
@@ 3.4 Tidal disruption of solitons
The tidal radius sets an upper limit on the size $\ell$ of the body, that is stable under tidal deformations. When $\ell>r_{\text{\tiny tid}}$, the body is susceptible to tidal disruption and can fragment into a collection of debris. [^24] While tidal disruption may not occur immediately when $\ell\lesssim r_{\text{\tiny tid}}$, it is important to note that tidal deformations can still take place. Indeed, in cases of significant deformation, after several cycles, the orbiting probe can become stretched to the extent that its size along the stretched direction approaches that of the tidal radius. This can ultimately result in the shattering of the body. This effect has been studied in fuzzy DM solitons orbiting in a NFW profile [181].
@@
The expected final picture is that the halo density distribution features a central superfluid core surrounded by superfluid debris left behind by the tidal disruption of outer solitons. These debris, when coarse-grained over distances greater than kpc, should approximately follow a NFW profile. Figure 3 provides a schematic representation of the thermal region of size $R_{\text{\tiny T}}$ following these events. However, since the above analysis relies on analytical estimates, numerical studies are clearly necessary to validate this conceptual framework or reveal its limitations. This would allow a deeper understanding of the complex non-linear dynamics at play in SDM structure formation. Recent works along this direction have already made some progress to understand the formation of SDM solitons in galaxy halos. For instance, through a 3D Schrödinger-Poisson simulation, Ref. [182] demonstrated that halos with sizes much larger than the healing length are likely to harbor the formation of a central soliton.
@@ 3.5 Summary
Dynamically, we expect halo formation to proceed as follows. In the first stages of its formation, the DM halo evolves in a similar way as standard collisionless DM. This continues until self-interactions start affecting the density profile, generating a region of the halo of size $R_{\text{\tiny T}}$ through thermalization, within which superfluiduity is established.
@@
Figure 4 shows a comprehensive plot of the parameter space of two-body interacting superfluids, for various values of the DM mass and cross section. The dotted, dashed and solid black lines correspond to solitons of size $\lambda_{\rm J}=0.1,2$ and $30$ kpc, respectively. The green region indicates the parameter space in which solitons are mainly sustained by quantum pressure, with density profiles resembling those of fuzzy DM models. The light purple region is excluded by the requirement that the cosmological pressure provided by the condensate is negligible at matter-radiation equality, which we will discuss in more details in Sec. 4.1. If we impose that condensates must be non-relativistic already at $T=\text{keV}$, then the excluded region is extended up to in between the dashed and solid black lines. The dark purple region corresponds to condensates that would have evaporated into dark radiation by the present time through number-changing processes (assuming real scalar condensates, since for a complex field the $U(1)$ symmetry protects the condensate from evaporating). Finally, the red stripe corresponds to the Bullet Cluster bound for degenerate DM particles, which will be discussed in detail in Sec. 5.1.
[…]
@@ 4.1 Early universe dynamics and bounds
DM particles in the sub-eV mass range must be produced as non-thermal relics, e.g., via the vacuum misalignment mechanism [185, 186, 187]. This mechanism, when applied to an ultra-light scalar field model, assumes that at some time in the early universe, the scalar field was initially configured homogeneously across the horizon scale. This configuration could have arisen from a prior phase transition, after which the scalar field acquired different values at different spatial locations, later homogenized by the Hubble flow.
How homogenization proceeds depends on the production time relative to cosmic inflation. If the phase transition occurs before inflation, the subsequent inflationary expansion naturally smooths out the field configuration. Conversely, if the transition takes place after inflation, the details of the smoothing process and the resulting homogeneous vacuum misalignment depend on the initial perturbation spectrum as well as on the non-linear dynamics following the phase transition, which are extensively studied through numerical simulations [188, 189, 190, 191, 192, 193]. In either case, the final outcome is a homogeneous classical field with a frozen amplitude.
@@
This can be somewhat problematic if our candidate makes up most of the DM, as it becomes relativistic too close to equality. Similarly, $\lambda_{\rm J}\gtrsim 10\,{\rm kpc}$ re-enter the cosmological horizon around the same time, and consequently fail to adequately seed structure formation (which would require an earlier horizon crossing) [194]. These results have also been confirmed through effective hydrodynamical simulations in Refs. [195, 196, 197], which showed that self-interacting scalar field DM models with core radii of order $1\,{\rm kpc}$ significantly suppress the formation of halos with masses below $10^{14}M_{\odot}$. Notably, this constraint is more stringent than the one estimated above. [^26] Similarly, Ref. [198] showed that the model could also be constrained using data from the cosmic microwave background, baryonic acoustic oscillations, growth factor measurements, and type Ia supernovae distances, resulting in a comparable bound on the size of the core.

### v2

@@ 2.6 Superfluidity and gravity
Before proceeding to the details, let us briefly comment on possible effects arising from quantum gravity. It is widely expected that gravity violates global symmetries: indeed, fundamental theories of gravity (such as string theory) appear incompatible with exact continuous global symmetries [39]. Hence, even if the starting point of the above discussion is a scalar field endowed with a fundamental $U(1)$ invariance, one should generically expect explicit $U(1)$-breaking mediated by operators suppressed by powers of the Planck scale.
However, as discussed in Sec. 2.4, the non-relativistic superfluid description remains valid even in the absence of a global $U(1)$ symmetry because the effective non-relativistic dynamics automatically restores particle-number conservation. Hence, gravity-mediated effects will merely generate effective quartic, sextic, and higher-order operators in the non-relativistic Lagrangian (2.33), suppressed by powers of the Planck mass. One can then expect to absorb this entire class of operators into a redefinition of the effective potential and, correspondingly, of the superfluid equation of state.
Substantial deviations would, however, arise when the relativistic modes of the condensate (or fast modes, in the terminology of [399, 400]) are taken into account, since relativistic corrections can induce a slow evaporation of the condensate, as we will discuss in Sec. 2.7.
@@ 2.6.1 Quasi-particle description
[…]
@@ 2.6.2 Hydrodynamical description
2.7 Quantum stability of superfluids and evaporation
The discussion above summarizes the semiclassical physics of non-relativistic superfluids, a regime well suited to describe SDM. A key element in constructing the EFT of superfluidity is the presence of a microscopic $U(1)$ symmetry that is spontaneously broken. However, non-relativistic condensates can also arise from a real scalar field theory, where the symmetry emerges in the non-relativistic limit. This occurs because the coarse-graining procedure over long time scales, employed to recover the Schrödinger equation from the real Klein-Gordon field (see e.g. [250] for this derivation in the context of fuzzy DM), effectively freezes the number-changing processes typical of theories without a conserved charge, leading to an emergent particle-number conservation. However, this emergent conservation is only approximate: number-changing interactions can appear through subleading secular corrections to the time-averaging procedure, ultimately causing the EFT of superfluidity for a real scalar to break down over sufficiently long timescales. In this Section, we quantify the magnitude of this effect.
To do so, we must examine the quantum dynamics of the superfluid. In particular, the exact time evolution of a quantum system is determined by the Hamiltonian dynamics of its corresponding quantum state. For coherent configurations, such as the condensates discussed above, the system can be represented, for example, by a coherent state or by a number eigenstate containing a large number of particles. A detailed treatment of these quantum states lies beyond the scope of this review. However, in Section 4.2, we provide a brief overview in the context of cosmological structure formation. Here, instead, we focus on outlining the key aspects of quantum time evolution that are relevant for such systems.
An important distinction arises depending on whether the underlying field is real or complex. In the latter case, one can show that superfluid configurations remain stable in the full quantum theory. The key point is that, because the relativistic Hamiltonian for the superfluid fluctuations (2.9) is time independent, the theory admits a vacuum that is exactly time-translation-invariant. This vacuum is precisely what one identifies with the superfluid state. The very existence of such a time-translation invariant vacuum for the fluctuations is what prevents a complex scalar field from decaying in time. Aspects of the quantum dynamics of complex scalar superfluids are discussed in Refs. [286, 357, 54] for the background evolution and stability, and in Refs. [15, 14, 411, 410] for the theory of fluctuations.
However, reiterating the derivation for a real scalar field shows that the theory of fluctuations becomes time independent only after applying the temporal smoothing procedure. Without this averaging, no time-translation invariant vacuum can be defined, and the superfluid fluctuations necessarily evolve nontrivially, inducing a corresponding nontrivial backreaction on the classical superfluid background. It is precisely this backreaction that drives the slow, gradual evaporation of the superfluid through number-changing processes among its internal constituents.
We now provide the depletion timescale of the real scalar condensate. This can be done by following the procedure outlined in [29, 58]. Let us consider the Heisenberg field equation of motion for a real scalar field with quartic interactions, where we consider a normalized potential $\frac{\lambda_{4}}{4!}\hat{\Phi}^{4}$. Next, we introduce the following field redefinition
$\hat{\Phi}(\vec{x},t)=v(t)+\hat{\psi}(\vec{x},t)\,,\qquad\langle\hat{\psi}(\vec{x},t)\rangle=0\,.$ (2.66)
Notice that, although it resembles the standard semiclassical expansion employed in the previous section, here the expansion is not organized around the classical solution to the equation of motion. By definition, here $v(t)$ is the one-point function of the field operator on the state that describes the condensate, and has to be determined. The time evolution of this function provides information about the evolution of the condensate. However, we have the freedom to choose the following initial conditions for this function
$\langle\hat{\Phi}(\vec{x},0)\rangle=v_{0}\,,\qquad\langle\partial_{t}\hat{\Phi}(\vec{x},0)\rangle=0\,$, (2.67)
which can be set by choosing the initial state of the system accordingly. To find the evolution of the one-point function, and hence of the quantum background, we need to solve the following system of differential equations
$(\partial_{t}^{2}+m^{2})v(t)+\frac{\lambda_{4}}{3!}v(t)^{3}=-v(t)\frac{\lambda_{4}}{2}\langle\psi^{2}(x)\rangle\,;$ (2.68)
$\left(\Box_{x}+m^{2}+\frac{\lambda_{4}}{2}v^{2}(t)\right)\langle\hat{\psi}(x)\hat{\psi}(y)\rangle=0\,.$ (2.69)
The derivation of this system, as well as its numerical solution, is provided in Refs. [29, 58].
This system is one-loop exact, and hence contains information on the leading non-trivial quantum correction to the dynamics of the condensate. The first equation represents the equation of motion for the background, corrected by one-loop quantum corrections on its right-hand side. Without these corrections, it would reduce to the standard classical equation of motion. The second equation represents the equation of motion for the two-point function of the fluctuation field. As we see, it is similar to the equation of motion of a scalar field in the vacuum, albeit corrected with a time and background-dependent mass. In order to solve this system, we use the initial conditions (2.67), and the fact that $\hat{\psi}$ behaves as a real scalar field with an effective mass $m_{\text{eff}}=m^{2}+\frac{\lambda_{4}}{2}v_{0}^{2}$ at time $t=0$.
The time evolution reveals that, while classically $v(t)$ is an oscillatory function describing a stationary configuration, at the one-loop level its amplitude is damped, reflecting the gradual loss of condensate particles through evaporation. The timescale at which the amplitude becomes half-damped (i.e., when an order-one fraction of the condensate has been depleted) matches the parametric estimate [172, 58]
$t^{-1}_{\text{dep}}\sim\lambda_{4}\left(\frac{\lambda_{4}\rho}{m^{4}}\right)^{3}m\,.$ (2.70)
This is the scale setting the survival time of the condensate.
Moreover, as a result of this evaporation, relativistic particles are produced through the conversion of the zero-momentum constituents of the condensate into a radiation component. It is important to emphasize that this particle production is fundamentally different from the standard parametric resonance mechanism described by the Mathieu equation [283, 317], which occurs only when the field lies within an instability band. By contrast, the process described above is generically present and does not rely on such conditions. Moreover, it is much slower, as it does not involve any exponential growth of mode functions. The standard exponential amplification can still occur, however, if the initial conditions are appropriately chosen [29].
@@ 3 Superfluid dark matter in galaxies
As shown above, the effect of gravity is balanced by two contributions. The first is the gradient of the polytropic equation of state, $\frac{1}{\rho}\vec{\nabla}P$, which encodes the role of self-interactions in the evolution of the system. The second is the quantum pressure contribution, $\frac{1}{2m^{2}}\vec{\nabla}\left(\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}\right)$, which arises from the wave-like nature of the light bosonic particles. Depending on which dominates at the characteristic scales of interest—typically $\sim{\rm kpc}$ scales in astrophysical contexts—different dynamical regimes emerge. If quantum pressure is the dominant contribution, the system behaves according to the fuzzy DM model. In the opposite limit, where self-interactions dominate, the dynamics are governed by the SDM model, which is the primary focus of this review.
@@ 3.1 Hydrodynamics of the dark matter superfluid
The first model of SDM was proposed by J. Goodman in the seminal paper [215]. In this work, the author introduced DM as a scalar boson with repulsive two-body interactions, which is degenerate at typical galactic densities and interacts through repulsive short-range interactions. The important implication of the original model is that the ground state pressure sourced by self-interactions can sustain self-gravitating, zero-temperature configurations. We refer to these structures as non-topological solitons. This stands in contrast to self-interacting DM models, where cores are prone to gravothermal collapse upon cooling (see the discussion at the end of Sec. 3.2[424, 35], which was one of the original motivations behind the introduction of the model.
@@
However, let us stress that early studies paid little attention to how the superfluid phase (and its associated solitonic cores) actually forms within DM halos. It was often assumed that once a halo reached equilibrium, it would settle into a simple configuration consisting of a single central soliton surrounded by an isothermal envelope [215]. In the next Section, we show that the reality is far more intricate. By following the evolution of a halo from an out-of-equilibrium state into the superfluid phase, we find that the interplay between DM thermalization and condensation leads to a much richer and more dynamic phenomenology than originally envisioned.
@@ 3.2 Formation of the superfluid phase
From these initial conditions, the system progressively thermalizes, approaching a maximum-entropy state in which a superfluid phase can develop. While the detailed evolution of these processes would need to be followed numerically, it is instructive to draw an analogy with the behavior of self-interacting DM. In this scenario, self-interactions drive the formation of approximately isothermal central regions, with the central cusp expanding and becoming shallower. However, this stage is only transient: as energy continues to flow outward, the core eventually undergoes gravothermal collapse, a runaway instability that would result in an even denser central region than the starting configuration.
This instability originates from the surprising thermodynamic fact that, for a self-gravitating system confined in a perfectly reflecting sphere, there is no global maximum of the entropy, but only local ones [21]. Moreover, if the confining sphere is allowed to expand beyond a critical radius, these local entropy maxima become inflection points, and the system becomes unstable. This process can be understood from the fact that virialized self-gravitating systems have negative heat capacity [320]. Hence, when a flow of heat is allowed from inner to outer regions of halos, such as in the presence of DM self-interactions, as the core loses energy, it shrinks and becomes hotter. If there is no way for this inner core to equilibrate with outer regions (that is, if the outskirts are quite insensitive to the heat transfer), the process continues until the system recollapses in a denser and compact core. The earlier phase of core expansion thus represents a transient stage during which the system attempts to reach thermal equilibrium. However, as the central distribution approaches a Maxwellian form and the system moves toward its local entropy maximum, the gravothermal instability sets in, driving the subsequent collapse.
If the system is assumed to maintain hydrostatic equilibrium and spherical symmetry [^21] See also Ref. [48] for a recent analysis on non‑spherical self-interacting DM halo shapes. throughout its evolution, the dynamics of the core collapse are governed by the gravothermal fluid equations, a coupled set of four differential equations [319, 184, 36, 282]. The first two equations enforce the hydrostatic equilibrium and the total mass conservation during all the stages of the collapse, and are provided by
$\frac{\partial M(r)}{\partial r}=4\pi r^{2}\rho(r)\,,\qquad\frac{\partial}{\partial r}\left[\rho(r)\langle v^{2}\rangle\right]=-\frac{GM(r)\rho(r)}{r^{2}}\,.$ (3.11)
Here, $\langle v^{2}\rangle$ is the local average squared velocity of particles in the halo which is determined in terms of a Maxwellian distribution, $M(r)$ is the mass enclosed in a radius $r$ from the center of the halo, and $\rho(r)$ is the local density. The second set of equations is provided by a heat and entropy equation
$\frac{L(r)}{4\pi r^{2}}=-\kappa\frac{\partial T(r)}{\partial r}\,,\qquad\frac{1}{{4\pi r^{2}}}\frac{\partial L(r)}{\partial r}=-\frac{\rho(r)\langle v^{2}\rangle}{\gamma-1}\left(\frac{\partial}{\partial t}\right)_{M}\log\left(\frac{\langle v^{2}\rangle}{\rho(r)^{\gamma-1}}\right)\,.$ (3.12)
Here, $T(r)$ is the local temperature, $L(r)$ is the energy flux, $\gamma$ is the adiabatic index of the DM fluid, and $\kappa$ is the conductivity coefficient. The Lagrangian time derivative $(\partial/\partial t)_{M}$ denotes differentiation at constant enclosed mass.
The conductivity coefficient $\kappa$ characterizes the efficiency of the system in transporting heat. It also plays a central role in connecting the microphysics of the fluid to the macroscopic transport behavior, since its value is determined by the way in which particles interact with one another. In models without self-interactions, the thermal conductivity is determined solely by gravitational interactions. While this contribution is negligible for DM halos, other systems, such as globular clusters, can experience significant heat transfer through gravitational encounters [320, 127, 325, 466]. When self-interactions dominate over gravitational effects in heat conduction, the thermal conductivity is approximately determined by [184, 459]
$\kappa\propto\frac{n\lambda^{2}_{\text{\tiny mfp}}}{t_{\text{\tiny rel}}}\,.$ (3.13)
Here, $n$ is the local DM density, $\lambda_{\text{\tiny mfp}}$ is the mean free path, which is connected to the underlying scattering cross section $\sigma$ by the relation $\lambda_{\text{\tiny mfp}}\sim 1/(\sigma n)$, while the relaxation time reads $t_{\text{\tiny rel}}\sim\lambda_{\text{\tiny mfp}}/v_{\text{\tiny rel}}$ in terms of the average relative velocity $v_{\text{\tiny rel}}$ of DM particles. Once the conductivity $\kappa$ is provided, one can solve Eqs. (3.11) and (3.12) and completely determine the dynamics of the gravothermal collapse for self-interacting DM halos.
Why does SDM behave differently in this respect? Unlike self-interacting DM, where the pressure from interaction is negligible on astrophysical scales, in SDM the high-density contracting core generates a pressure from particle interactions that counteracts the usual gravothermal contraction. From a thermodynamic perspective, this effectively reverses the sign of the heat capacity of the core. Consequently, the core cools, gradually re-equilibrates with the outskirts, and condenses. Collapse into an extremely dense configuration is avoided, and the system settles into the cored structures described in the previous section. As noted above, however, the formation of the superfluid phase has not yet been numerically validated in the literature.
@@ 3.4 Tidal disruption of solitons
The tidal radius sets an upper limit on the size $\ell$ of the body, that is stable under tidal deformations. When $\ell>r_{\text{\tiny tid}}$, the body is susceptible to tidal disruption and can fragment into a collection of debris. [^28] While tidal disruption may not occur immediately when $\ell\lesssim r_{\text{\tiny tid}}$, it is important to note that tidal deformations can still take place. Indeed, in cases of significant deformation, after several cycles, the orbiting probe can become stretched to the extent that its size along the stretched direction approaches that of the tidal radius. This can ultimately result in the shattering of the body. This effect has been studied in simulations of fuzzy DM solitons orbiting in a NFW profile [211], which showed that repulsive self-interactions accelerate disruption, and that there is a degeneracy between soliton mass and self-coupling strength in determining disruption timescales.
@@
The expected final picture is that the halo density distribution features a central superfluid core surrounded by superfluid debris left behind by the tidal disruption of outer solitons. These debris, when coarse-grained over distances greater than kpc, should approximately follow a NFW profile. Figure 3 provides a schematic representation of the thermal region of size $R_{\text{\tiny T}}$ following these events. However, since the above analysis relies on analytical estimates, numerical studies are clearly necessary to validate this conceptual framework or reveal its limitations. This would allow a deeper understanding of the complex non-linear dynamics at play in SDM structure formation. Recent works along this direction have already made some progress to understand the formation of SDM solitons in galaxy halos. For instance, through a 3D Schrödinger-Poisson simulation of a virialized halo plus stochastic fluctuations in repulsive scalar‐field DM models, Ref. [203] demonstrated that halos with sizes much larger than the healing length are likely to harbor the formation of a central soliton, with halo assembly history potentially leading to a large scatter in the soliton masses in cosmological halos.
@@ 3.5 Summary
Dynamically, we expect halo formation to proceed as follows. In the first stages of its formation, the DM halo evolves in a similar way as standard collisionless DM. This continues until self-interactions start affecting the density profile, generating a region of the halo of size $R_{\text{\tiny T}}$ through thermalization, within which superfluidity is established.
@@
Figure 4 shows a comprehensive plot of the parameter space of two-body interacting superfluids, for various values of the DM mass and cross section. The dotted, dashed and solid black lines correspond to solitons of size $\lambda_{\rm J}=0.1,2$ and $30$ kpc, respectively. The green region indicates the parameter space in which solitons are mainly sustained by quantum pressure, with density profiles resembling those of fuzzy DM models. The light purple region is excluded by the requirement that the cosmological pressure provided by the condensate is negligible at matter-radiation equality, which we will discuss in more details in Sec. 4.1. If we impose that condensates must be non-relativistic already at $T=\text{keV}$, then the excluded region is extended up to in between the dashed and solid black lines. The dark purple region corresponds to condensates that would have evaporated into dark radiation by the present time through number-changing processes, assuming that DM is comprised of real scalar bosons. In particular, within this range of parameters, one would find that the typical depletion time-scale of the condensate (2.70) would be shorter than a Hubble time. Finally, the red stripe corresponds to the Bullet Cluster bound for degenerate DM particles, which will be discussed in detail in Sec. 5.1.
We identify two allowed sub-regions in the plot. The white region on the left of the red stripe corresponds to halos in which only central regions present a superfluid phase. In this case, $R_{\text{\tiny T}}<R_{\text{\tiny V}}$, allowing only solitons up to $\lambda_{\rm J}\sim 6$ kpc. Conversely, the white region on the right side corresponds to halos that are entirely in thermal equilibrium ($R_{\text{\tiny T}}>R_{\text{\tiny V}}$). In this case, Jeans scales up to $\mathcal{O}(100\text{ kpc})$ are allowed, provided that the Bullet Cluster bound can be relaxed for halos in global thermal equilibrium. However, as we will show in the next Section (see also the light purple region), these values are marginally consistent with having a non-relativistic equation of state at matter-radiation equality.
@@ 4.1 Early universe dynamics and bounds
DM particles in the sub-eV mass range could be produced as non-thermal relics, [^30] Sub-eV mass bosons can also be produced thermally in the early universe through interactions with SM particles. However, such a component would contribute as dark radiation, thus being constrained by measurements of the effective number of relativistic species through CMB and BBN data [107, 108, 339, 220], or suppress structure formation through free streaming effects [458]. e.g., via the vacuum misalignment mechanism [381, 2, 158], to make up a significant fraction of the DM in the universe. This mechanism, when applied to an ultra-light scalar field model, assumes that at some time in the early universe, the scalar field was initially configured homogeneously across the horizon scale. This configuration could have arisen from a prior phase transition, after which the scalar field acquired different values at different spatial locations, later homogenized by the Hubble flow.
How homogenization proceeds depends on the production time relative to cosmic inflation. If the phase transition occurs before inflation, the subsequent inflationary expansion naturally smooths out the field configuration. Conversely, if the transition takes place after inflation, the field takes random values in causally disconnected regions, leading to large inhomogeneities and the formation of topological defects [275, 46]. In this post-inflationary case, the spontaneous breaking of the global UU(1) symmetry triggers, via the Kibble–Zurek mechanism, the formation of a vortex network with an initial coherence length set by the Hubble radius. As the horizon grows, regions containing vortices come into causal contact, temporarily increasing the apparent inhomogeneity. However, analytical and numerical studies of global-string networks show that their non-linear evolution rapidly drives the system toward a scaling regime with roughly one string per Hubble patch [233, 269, 445, 106, 359, 360]. In this attractor state, vortices intercommute, self-intersect, and fragment into loops that efficiently radiate the Goldstone mode, converting the network’s energy into phonons on Hubble timescales. Provided the transition occurs well before matter-radiation equality, the decay of loops and cosmic expansion wash out large-scale gradients, yielding an effectively homogeneous superfluid background on Hubble scales.
@@
This can be somewhat problematic if the candidate makes up most of the DM, as it becomes relativistic too close to equality. Similarly, $\lambda_{\rm J}\gtrsim 10\,{\rm kpc}$ re-enter the cosmological horizon around the same time, and consequently fail to adequately seed structure formation (which would require an earlier horizon crossing) [160]. These results have also been confirmed through effective hydrodynamical simulations in Refs. [138, 408, 228], which showed that self-interacting scalar field DM models with core radii of order $1\,{\rm kpc}$ significantly suppress the formation of halos with masses below $10^{14}M_{\odot}$. Notably, this constraint is more stringent than the one estimated above. [^31] Similarly, Ref. [227] showed that the model could also be constrained using data from the cosmic microwave background, baryonic acoustic oscillations, growth factor measurements, and type Ia supernovae distances, resulting in a comparable bound on the size of the core.